OpenStudy (anonymous):

a particle of unit mass falls from rest, under gravity, in a medium in which there is a resistance of magnitude gv^2/U^2, wher v is the speed of the particle after time t and u is a positive constant. show that 2gt=U ln(U+v/U-v).

OpenStudy (anonymous):

Apply Newtons 2nd law. \[m(d ^{2}x/dt ^{2}) = mg-V ^{2}/U ^{2}\] we also know \[d ^{2}x/dt ^{2} = dV/dt\] set m=1 substitute into newtons 2nd law and integrate (hint use a trigonometric substitution . change of variable for V)